OpenStudy (anonymous):

a train accelerates from rest at 1.00m/s^2. at the same instant as the 1st train starts, another train is 1000m away from it and moving towards it at a constant speed of 20m/s. how far from the starting point of the 1st train are the 2 when they meet? please use UAM formulas :D

OpenStudy (anonymous):

V=speed x=distance t=time a=acceleration V=x/t a=V/t so x.t equal V/t t^2=x/a so "t" is a same point of this question. u should use t if 2 trains meet x point first train came "x" distance, 2. one came "1000-x" km t=V/x same time t^2=x/a now u should fill in the blanks in this problem ;) i cant write here all of this equation.

OpenStudy (anonymous):

first train distance is x second train distance is 1000-x

OpenStudy (anonymous):

please explain further. :(